Two numbers differ by 6. Five times the smaller is 6 more than three times the larger. What are the numbers?
Answer by mananth(16949)

if one number is x
the other will be x+6
...
5x=3(x+6)+6
5x=3x+18+6
5x-3x=24
2x=24
/2
x= 12
...
the other number is x+6 = 12+6 =18
...
m.ananth@hotmail.ca